OpenStudy (anonymous):

d^2y/dx^2=9y solve for y

OpenStudy (anonymous):

let \(\huge \frac{dy}{dx}=p \) and note that \(\huge \frac{dp}{dx}=\frac{dp}{dy}\frac{dy}{dx}=\frac{dp}{dy} p \)

OpenStudy (anonymous):

i know that the answer is e^(-3x) or e^(3x) im just not sure about the steps

OpenStudy (unklerhaukus):

now simply separate the variables \[\frac{\text d^2y}{\text dx^2}=9y\] \[\frac{\text dp}{\text dy}p=9y\] \[p{\text dp}=9y\text dy\] and integrate \[\int p\cdot{\text dp}=9\int y\cdot\text dy\]

OpenStudy (unklerhaukus):

\[\frac{p^2}2=9\frac{y^2}2+c_1\]\[{p^2}=9y^2+c_2\] \[p=\sqrt{9y^2+c_2}\] \[\frac{\text dy}{\text dx}=\sqrt{9y^2+c_2}\] \[\frac{\text dy}{\sqrt{9y^2+c_2}}=\text dx\] \[\frac 13\int\frac{\text dy}{\sqrt{y^2+c_2}}=\int\text dx\]

OpenStudy (unklerhaukus):

that should be \(c_3\) in the last equation, \(c_2/9=c_3\)

OpenStudy (anonymous):

thanks!

OpenStudy (unklerhaukus):

There is an easier method;\[\frac{\text d^2y}{\text dx^2}=9y\]\[\frac{\text d^2y}{\text dx^2}-9y=0\]\[m^2-9=0\]\[(m+3)(m-3)=0\]\[m=\pm 3\]

OpenStudy (anonymous):

Thanks that is easier to follow

OpenStudy (anonymous):

If it was -9y rather than 9y would your only option be to use the first method

OpenStudy (unklerhaukus):

\[\frac{\text d^2y}{\text dx^2}=-9y\]\[\frac{\text d^2y}{\text dx^2}+9y=0\]\[m^2+9=0\]\[(m+3i)(m-3i)=0\]\[m=\pm 3i\]

OpenStudy (anonymous):

which correlates to sin3x or cos3x?

OpenStudy (unklerhaukus):

\[\frac{\text d^2y}{\text dx^2}=9y\qquad\Longrightarrow\qquad y(x)=Ae^{3x}+B^{-3x}\] \[\frac{\text d^2y}{\text dx^2}=-9y\qquad\Longrightarrow\qquad y(x)=A\cos{3x}+\sin{3x}\]

OpenStudy (unklerhaukus):

ops there should be a constant on the last term of the last line \(\cdots+B\sin3x\)

OpenStudy (anonymous):

great! thanks again!